OpenStudy (anonymous):

solve for x: ln(5x+1)+lnx=ln4

OpenStudy (callisto):

ln(5x+1)+lnx=ln4 ; ln(5x+1)(x) = ln4 ; 5x^2 + x- 4 =0 ; x = 1 or x=-0.8 (rejected)

OpenStudy (callisto):

sorry , x=0.8 or x=-1(rejected)

OpenStudy (callisto):

Once again, ln(5x+1)+lnx=ln4 ; ln(5x+1)(x) = ln4 ; 5x^2 + x- 4 =0 ; x=0.8 or x=-1(rejected)

OpenStudy (anonymous):

thanks!